OpenStudy (anonymous):

A hiker in Africa discovers a skull that contains 63% of its original amount of C-14. N=Noe^-kt No=initial amount of C-14(at time t=0) N=amount of C-14 at time t t=time, in years k=0.0001

OpenStudy (anonymous):

Ok so we have: \[ N(t)=N_o e^{-kt}\]

OpenStudy (anonymous):

We are told that their is 63% of C14 remaining in the material hence: \[ \frac{N(t)}{N_o}= 63\% =0.63= e^{-kt}\]

OpenStudy (anonymous):

yes but it says find the age of the skull to the nearest year.

OpenStudy (anonymous):

Taking the ln of both sides: \[\ln(0.63)= -(0.0001 yr^{-1})t\]

OpenStudy (anonymous):

Divide and you have you answer

OpenStudy (anonymous):

divide 0.63 / -0.0001?

OpenStudy (anonymous):

Please note it isnt 0.63 but ln(0.63)

OpenStudy (anonymous):

And then it is a yes

OpenStudy (anonymous):

Btw, in case your worried about the minus sign... the fact that the argument of the logarithm is <1 means its result will be <0.... so the minus signs cancel and it will yield a positive value for t

OpenStudy (anonymous):

i got -6300

OpenStudy (anonymous):

natrual log of 0.63 divided by -10^-4

OpenStudy (anonymous):

Please note that the CRUCIAL step in this problem was taking the log of both sides. Without that step you cannot reduce the exponential factor.

OpenStudy (anonymous):

your confusing me

OpenStudy (anonymous):

What do you find confusing?

OpenStudy (anonymous):

Please, I cant help if I dont know where your getting confused.

OpenStudy (anonymous):

Go through the process I did at the top and let me know which step you are uneasy about and I will give a more detailed explanation.

OpenStudy (alekos):

show him the steps in detail

OpenStudy (anonymous):

I was under the impression I did alekos but I will go through it again.

OpenStudy (anonymous):

Ok so we given in the problem that the amount of C14 we find in the skull is 63%

OpenStudy (anonymous):

This means that of the original amount: \[N_0=100\% (some \ amount \ of \ nuclei)=(some \ amount \ of \ nuclei) \] It decays according to the law \[N(t)=N_0 e^{-kt}\] Where the N's correspond to the number of nuclei of the material in quesion remaining (N(t)) or the original amount (N_0) So that at time T, we have: \[N(T)=63\% (some \ amount \ of \ nuclei)=(some \ amount \ of \ nuclei) e^{-kT}\] Now whatever the total amount of nucleai there are in the sample is irrelevant because I can just divide both sides by that amount leaving just the percent remaining.

OpenStudy (anonymous):

This results in the equation I gave above: \[63\% = 0.63=e^{-kT}\]

OpenStudy (anonymous):

Now just like when we have an equation of the form: \[2=x^2\] Which we solve by applying the inverse function of squared (i.e. the square root) to both sides (ignoring the +/- that comes in for simplicity): \[\sqrt{2}=\sqrt{x^2}=x\] When we have an exponential equation of the form: \[2=e^x\] We apply the inverse function to the exponential function (aka the natural logarithm) to both sides in order to simplfy and solve: \[ \ln(2)=\ln(e^x)=x\]

OpenStudy (anonymous):

Note that an exponential function can NEVER be negative (unless I explicitly multiply it by a -1), which means when I do this inverse function business applying the logarithm, I dont have to worry about a +/- like you do when you apply a square root. So don't let that trouble you.

OpenStudy (anonymous):

So performing this step (taking the natural log of both sides of the equation) yields: \[\ln{0.63}=\ln{e^{-kT}}=-kT=-(10^{-4}yr^{-1})T\] Then solve for T, by dividing both sides by k: \[T=\frac{\ln{0.63}}{-(10^{-4}yr^{-1})}\]

OpenStudy (anonymous):

Please calculate this value and express it with the appropriate units here I and I will check your answer.

OpenStudy (anonymous):

@ix.ty are you still here?

OpenStudy (anonymous):

im on a diffrent question now

OpenStudy (anonymous):

A country's population in 1992 was 222 million. In 2001 it was 224 million. Estimate the population in 2004 using the exponential growth formula. Round your answer to the nearest million. P = Aekt

OpenStudy (anonymous):

So what did you get for the last one?

OpenStudy (anonymous):

i got it wrong that was my last question but i passed the test

OpenStudy (anonymous):

Oh so before I even try to start what looks to be a HARDER question on the same material it bears going over the last problem to iron out what confused you. Please tell me what you found confusing.

OpenStudy (anonymous):

No judgement, I fancy myself like a bit of an auto-mechanic.... I need to know where the problem is in order to try and fix it.

OpenStudy (alekos):

dont bother plasma, seems like you're wasting you're time

OpenStudy (anonymous):

Perhaps, but I am stubborn and banging my head against the wall might be a bit of a pastime for me.... But either way I am willing to help @ix.ty but I will only stick around for maybe 10-20 minutes (Ill check back at this tab) then I will move on